In a isosceles triangle ABC, b is the vertex. the vertex angle measures 3 times the sum of the measures of the other 2 angles. what is the measure of angle b?
is it 36
is it 72
is it 108
is it 135
i need help soon as possible

Answer by checkley77(12844)

b=3*2a
b+a+a=180
3*2a+2a=180
6a+2a=180
8a=180
a=180/8
a=22.5 degrees for the other 2 angles.
b+2*22.5=180
b+45=180
b=180-45
b=135 degres for the apex angle.